### v2

# Verifying Device Drivers with Pancake

Abstract
Device driver bugs are the leading cause of OS compromises, and their formal verification is therefore highly desirable. To the best of our knowledge, no realistic and performant driver has been verified for a non-trivial device. We propose Pancake, an imperative language for systems programming that features a well-defined and verification-friendly semantics. Leveraging the verified compiler backend of the CakeML functional language, we develop a compiler for Pancake that guarantees that the binary retains the semantics of the source code. Using automatic translation of Pancake to the Viper SMT front-end, we verify a performant driver for an Ethernet NIC.
1 Introduction
Device driver bugs are the leading cause of OS compromises, accounting for the majority of the 1,057 CVEs reported for Linux in the period 2018–22 [MITRE Corporation, 2023]—clearly they should be the #1 targets of OS verification efforts.
While there have been prior efforts to verify drivers [Alkassar and Hillebrand, 2008; Alkassar, 2009; Duan and Regehr, 2010; Duan, 2013; Chen et al., 2016; Kim et al., 2008; Penninckx et al., 2012; Möre, 2021], to our knowledge none have yet succeeded on realistic, non-trivial devices, nor have they presented any performance analysis of the drivers verified.
Most of these efforts used highly manual interactive theorem proving, and sometimes required drivers to be written and analysed in assembly. On the other hand, attempts to apply more usable methods like model checking [Kim et al., 2008] and automated deductive verification [Penninckx et al., 2012] left significant gaps between the analysed model and the real code.
Of the above, only Chen et al. [2016] provided an end-to-end verification story that preserved the driver’s correctness from a driver-appropriate systems programming language down to the binary. Likewise, with no formal semantics or verified compiler, recent proposals like that of [Chen et al., 2024] to verify drivers written in Rust [Klabnik and Nichols, 2017] have no plan to close the semantic gap to the binary.
Devices are commodities: new ones are created all the time. Therefore, any practicable driver verification must have a high degree of usability and automation. Drivers are typically written in C, but verification of C code is made needlessly expensive by C’s complicated semantics. In short, the situation demands a performant systems language with support for usable automated verification, and a means of ensuring the properties of verified drivers are preserved at the binary level.
In this paper, we present verification of a performant, real-world Ethernet driver that demonstrates a new approach to efficient development of performant and verified device drivers, achieved through the choice of driver design and the use of Pancake, a new imperative programming language designed specifically for verification of low-level systems code. Namely, we verify the Pancake version of an Ethernet driver for LionsOS [Heiser et al., 2025], which uses single-threaded, single-purpose (“modular”) driver designs as advocated by Ryzhyk et al. [2009a, 2010] and outperforms Linux networking with its standard C driver. The Pancake version of the driver we verify also enjoys the excellent performance owing to the simple design, with reasonable overhead coming from porting to Pancake. It will be made available as a choice of network drivers for LionsOS soon.
Our new programming language Pancake has several features that facilitate verification, the most significant being:
  • a verified compiler from Pancake to binary that leverages the final stages of the verified CakeML compiler [Tan et al., 2019]. This guarantees that the semantics of drivers and other systems code written in Pancake, will be preserved at the binary level. Section 2 discusses the overall structure of Pancake’s semantics and the compiler.
  • an automated deductive verification front-end for Pancake that leverages the Viper verification framework [Müller et al., 2016b], a middle-end for various SMT solver-based verification back-ends. This takes the form of (1) an annotation syntax for Pancake, and (2) a transpiler from annotated Pancake to the Viper intermediate language (IL), which we explain in Section 3.
Using this support, we produce the first verification of a device driver for a non-trivial device, a driver for a popular 1 Gb/s Ethernet card for the seL4-based LionsOS [Heiser et al., 2025]. This paper reports on the design of the Pancake language, mainly on its compiler and semantics, and on the verification of the Pancake Ethernet driver using the Viper front-end for Pancake.
Our experience shows that Pancake’s automated deductive verification support is easy to use for those with a systems development background, allowing the verification of critical guarantees for practical drivers (Section 4). The verification of our Ethernet driver took around three person-months. The driver was ported from C to Pancake and then verified by the one member of the team who had no prior verification experience. We expect the process will be even faster once a library of example verified drivers is available. This demonstrates Pancake as a viable alternative to C for systems-level code, but one with the advantage of accessible end-to-end verification support.
Our evaluation in Section 5 shows that the verified Pancake driver performs very close to the C version.
In Section 6 we discuss the trusted computing base (TCB) and threats to the validity of Pancake’s automated deductive verification story and end-to-end semantic preservation, and our plan to address them with further verification.
2 Pancake Language and Compiler
2.1 The Rationale
[…]
Many attempts to make better systems programming languages incorporate advanced language features to make safety properties hold by construction. For example, Cogent’s linear type discipline prevents memory leaks [Amani et al., 2016], Rust’s borrow checker enforces ownership and lifetimes [Klabnik and Nichols, 2017], and Cyclone incorporates garbage collection and ML-style polymorphism [Jim et al., 2002]. Such features can eliminate whole classes of bugs, or at least reduce bug density, but at the cost of complicating the language semantics and implementation. Garbage collection introduces unpredictable delays that are highly undesirable in low-level systems code. Yet these approaches still fall short of ensuring full functional correctness, and it is unclear how helpful they are in achieving it.
Functional correctness proofs routinely rely on significantly stronger properties than type systems typically guarantee. A stronger type system can give more properties, but unless it is so powerful (and undecidable) as to be a full-featured proof calculus, functional correctness proofs will need more.
The information provided by a type system is only useful in verification if the type system is sound, but most practical languages have unverified type systems, or type systems with known soundness bugs. Type systems can be verified [Naraschewski and Nipkow, 1999], but type soundness proofs are delicate, and have subtle interactions with even minor language changes. Maintaining a type soundness proof for a living language can significantly bog down development.
Moreover, the safety guarantees of a language only hold if no backdoors are used. But low-level systems programs often need to break out of type-safe environments. For example, driver code must adhere to hardware-specified data locations, layouts and access protocols. Hence driver code in safe languages must use significant amounts of unsafe code, effectively escapes to C [Astrauskas et al., 2020; Evans et al., 2020], which mostly eliminates the benefits of safe languages.
Instead of adding more safety features to a language, which tends to complicate the semantics, we believe simple proofs require a simple formal semantics. Of course, the semantics must exist in the first place: despite years of research [Jung et al., 2018; Weiss et al., 2019; Kan et al., 2018; Wang et al., 2018], there is still no complete formal specification of Rust.
We propose Pancake as the solution—a radically minimal language that nonetheless offers a sufficiently expressive interface for writing low-level systems programs, such as device drivers, alongside a number of advantages for formal verification. Most importantly, the language is completely specified by a straightforward formal semantics that fits in a few hundred lines of HOL4 code, with a simple memory model, no notion of undefined behaviour, and no ambiguities in evaluation order.
Pancake is an unmanaged language with no static type system, at a level of abstraction between C and assembly. The data representation and memory models are kept as simple as possible. The only kinds of data are machine words, code pointers, and structs. Programs cannot inspect the stack, which simplifies semantics. All memory is statically allocated; there is no equivalent of malloc and free.
2.1.1 Non-Goals
Concurrency
User-level drivers running as separate processes do not need to be multi-threaded. Even for Linux in-kernel drivers, Ryzhyk et al. [2009a, 2010] demonstrate that single-threaded drivers are feasible and performant. They furthermore show that 19% of Linux driver bugs are concurrency bugs, which are automatically eliminated by such a design. Single-threaded drivers are much easier to verify, and map well onto the modular design of microkernel-based OSes. This approach is routinely used for drivers on seL4, including in LionsOS, which outperforms Linux on networking using this design [Heiser et al., 2025].
Therefore, Pancake has no concurrency primitives, and our Ethernet driver has no internal concurrency. We will see that this does not meaningfully inhibit performance.
Rust-level verification
Pancake is not intended to replace Rust; rather, it targets a different niche.
Rust’s safety features benefit developers by allowing them to write more trustworthy code in the absence of verification, and Rust-based verification tools like Verus and Prusti have been very successful. However, Rust is a complex language with no authoritative formal semantics, putting any verification on soft foundations. It also requires trusting a large, unverified compiler, which has been found to yield memory-unsafe binaries from typechecked memory-safe Rust code [Speykious, 2024]. Furthermore, “unsafe” code is required for implementing device drivers in Rust, and existing tools have limited support for verification of unsafe Rust.
This level of verification may be sufficient for certain situations, but we aim for stronger guarantees. We take the compiler out of our TCB by fully verifying it, based on Pancake’s precise semantics. A simpler language with formal semantics makes verification easier in both automated and interactive settings. Overall, what we offer is a framework for producing secure low-level components with a very small TCB, through Pancake’s formal semantics and verified compiler.
2 fun example(1 N, 1 shared_addr) {
3 /@ requires mem_access(N) @/
4 /@ ensures mem_access(N) @/
5 /@ ensures N == old(N) @/
6 /@ ensures shared_addr == old(shared_addr) @/
7 var x = 0;
8 var i = 0;
9 while (true) {
10 /@ invariant 0 <= i && i < N && x <= 42 @/
11 /@ invariant mem_access(N) @/
12 var new_x = lds 1 @base + @biw * i;
13 if (i + 1 == N) {break;}
14 if (new_x > 42) {break;}
15 x = new_x;
16 i = i + 1;
17 }
18 /@ assert x <= 42 @/
19 !stw shared_addr, x;
20 return 0;
21}
Listing 1: Pancake code snippet (concrete syntax) with annotations.
2.2 The Language and its Semantics
Pancake looks and feels like a traditional imperative language (see 1, ignore the /@...@/ annotations for now). Figure 1 shows the current abstract syntax of Pancake. The nuts and bolts should be familiar to any programmer. Mutable variables, if, while—nothing fancy. This is a deliberate design decision: we want Pancake to feel simple and familiar to systems programmers. In our experience, this has been borne out in practice: systems programmers familiar with C have found Pancake easy to learn.
[…]
Figure 1: Abstract syntax of Pancake.
Another key design concern is to give programmers direct access to low-level details without the language getting in their way. This is part of the motivation for Pancake’s perhaps most radical design decision: no static type system, and no distinction between different kinds of data. In sloganeering terms, Pancake is a language where everything is a machine word. For example, there is no distinction between pointers and integers: it’s all words.
We can treat a word as an integer by adding to it, or treat it as a pointer by dereferencing it. The programmer can do pointer arithmetic freely, which is of course unsafe in general. We do not attempt to make it safe; rather, we give it a simple and well-defined semantics that can support formal verification, without the need for complicated rules about (say) pointer provenance. In this way, the historically minded reader may find the language design closer to BCPL than C.
With this in mind, the data representation and memory models of Pancake are kept as simple as possible. There are only three kinds of data: machine words, code pointers, and structs (whose fields are machine words, code pointers, or nested structs). Local variables are stack-allocated, and we do not allow pointers into the stack. Global data may be stored in a statically allocated global memory region.
The operational semantics of Pancake is specified using functional big-step semantics [Owens et al., 2016]. In this style, the core of the semantics is a logical function that looks like an interpreter, but is not necessarily executable. This simplifies formal proofs of compiler correctness by making the semantics more amenable to term rewriting. The semantics of a program is defined in terms of how it communicates with the outside world; specifically, as a possibly infinite trace of I/O events, each of which denotes either a foreign function call or a shared memory load/store operation. The semantics is parameterised on a model of the outside world and the kind of I/O events that can occur.
2.3 Verified Compiler
Figure 2: CakeML and Pancake compiler stack.
The Pancake compiler is verified, and thus not in the TCB. This sidesteps the need for fragile validation of compiler output on a program-by-program basis [Sewell et al., 2013].
We reuse the lower parts of the verified CakeML compiler [Kumar et al., 2014; Tan et al., 2019], which compiles CakeML programs to machine code for six architectures. Figure 2 shows the CakeML compiler on the right and the Pancake compiler on the left. Arrows denote compilation passes, and yellow boxes denote intermediate representations.
CakeML itself is a high-level functional programming language, unsuited for low-level systems programming. For example, CakeML’s memory management is all handled by the language runtime, and memory allocation may trigger a stop-the-world garbage collection at any time.
By contrast, Pancake is explicitly designed to be unmanaged and close to hardware, and to have no runtime. Yet by integration into the CakeML ecosystem, it can reuse many of the existing correctness proofs for the CakeML compiler. The CakeML compiler provides backends for various target architectures [Tan et al., 2019] verified against detailed models of the target ISA semantics.
The first few phases of the Pancake compiler go through two new intermediate languages (ILs). The first compiler phase flattens structs and converts the programs to CrepLang, which is a stepping stone into LoopLang. In LoopLang, we compute minimal live sets and divide loops (including their break and continue statements) into tail-recursive functions that better fit the CakeML IL called WordLang. When the program under compilation is translated from LoopLang to WordLang, all loops are replaced with fast tail-calls, as WordLang has no loops. Later, as in CakeML, these tail calls will be realised as fast simple jumps in the generated machine code.
Once we have entered WordLang, we use the CakeML compiler’s compilation phases, including its phases for instruction selection, register allocation, concretisation of the stack, and, at the end, encoding of the entire program into concrete machine code.
The compiler correctness proofs allow safety and liveness properties of Pancake programs to carry over to the machine code that runs them.
Our Pancake compiler has a special feature which produces not only the output machine code, but also predicts the concrete maximum stack size the program can use, as long as the input program has recursive calls only in tail position. We have proved that this bound suffices for running the Pancake program without running out of stack space. This allows Pancake to guarantee the absence of premature termination arising from stack overflows.
The functional correctness proof of the Pancake compiler follows the style of the CakeML proofs. The top-level correctness theorem states that the source Pancake program and the compiler-generated machine code exhibit exactly the same observable I/O events, if the compiler-generated code is run with enough stack space.
The compiler and the associated proofs are open source. [^1] http://code.cakeml.org for source, or http://cakeml.org for pre-packaged versions. Note that Pancake is fully integrated into the CakeML compiler, rather than a stand-alone release.
2.4 Challenges with Device Drivers
Using CakeML as Pancake’s backend compiler presents a number of issues when compiling device driver code. First, CakeML assumes exclusive ownership of its statically allocated memory region, and that this memory is not observable by the environment. These assumptions are not valid for memory-mapped device registers, which are special memory locations used for interfacing with device hardware, nor do they hold when devices directly write to memory (DMA). They also do not fit the LionsOS model, which heavily depends on shared-memory communication between OS components [Heiser et al., 2025].
With this limitation, any interaction with the device from CakeML would have to be mediated by FFI calls to C, adding a layer of indirection to the TCB. We eliminate this indirection in Pancake by adding native support for shared memory.
[…]
We add shared memory semantics to all compiler ILs from Pancake downwards, and verify the full compilation chain. We give load/store operations on shared memory regions a special syntax (ShMemLoad and ShMemStore in Figure 1). In the semantics, these operations are considered observable events. At the end of the compiler pipeline they are converted to ordinary loads and stores.
Another challenge arising from the CakeML lineage is the issue of function entry points. The CakeML compiler expects code to only ever be called via the main function, and to exit entirely when main returns. Moreover, main is preceded by lengthy initialisation, and does not support parameters.
Device drivers in LionsOS must support several functions as entry points, and for these to support re-entry. Thus, using CakeML’s execution flow demands workarounds, such as branching in the main function based on data indirectly passed through memory, and manually editing the compiler output to return to the caller. Even then, a substantial performance penalty arises from repeated re-initialisation. This is clearly neither scalable nor performant.
We address this by implementing support for multiple entry points. A function declaration can be tagged with the export keyword, which extends the generated compiler output to expose the function to the calling conventions of the target platform, and to restore the state initialised by the main function when the exported function is called. This circumvents the re-initialisation, and handles argument passing and returning to the caller without programmer intervention. It enables calling into the Pancake driver as if it were written in C. The top-level compiler correctness proof currently accounts only for the main entry point in that some of the assumptions are too specific to initialisation, which need to be relaxed to account for the re-entry points.
2.5 Verification Approaches
For the purpose of using Pancake to produce high-assurance device drivers, the verified compiler (Section 2.3) is only half of the story: we must also be sure that the Pancake code implements the device driver correctly (incorrect source $+$ verified compiler $=$ incorrect machine code).
The gold-standard (but potentially expensive) way to verify source code is to use interactive theorem-proving (ITP), as done by (for example) the seL4 [Klein et al., 2014] and CertiKOS [Gu et al., 2016] kernels, and the CompCert [Leroy, 2009], and CakeML [Kumar et al., 2014] (and thus Pancake) compilers. As Pancake is an imperative language, it would be natural to use an ITP-based implementation of Hoare logic. We describe preliminary and future work on ITP-based approaches to source verification in Section 6.2.
For productivity, given the commodity nature of devices, we use automated deductive verification, rather than an ITP. Using Viper [Müller et al., 2016b] (and its realisation of Hoare logic) increases the size of the trusted computing base (see Section 6.1), but can be expected to be less expensive in terms of developer time and expertise. Our case study bears this out.
To support this, Pancake syntax includes Viper annotations (visible in 1). They resemble comments in the source code, and can appear at the top level and within function bodies. Top-level annotations are used, for instance, to specify function contracts, while in-function annotations can be used for loop invariants. The Pancake compiler ignores annotations, but the Viper transpiler (Section 3) preserves the annotations so that verification can be carried out.
3 A Viper Front-end for Pancake
Our verifier transpiles annotated Pancake code into the Viper IL, then verifies the generated Viper code using Viper’s symbolic execution backend Silicon [Schwerhoff, 2016]. This approach is comparable to that of Viper front-ends such as Gobra [Wolf et al., 2021] for Go and Prusti [Astrauskas et al., 2022] for Rust.
Our transpiler first uses a diagnostic feature of the Pancake compiler to extract the abstract syntax tree (AST) of the input program, which includes Pancake-level (1) annotations of the kind shown in 1 as well as (2) hooks into a device model that we will describe in Section 3.3. It then translates that AST into our chosen encodings for Pancake’s variables, values, memory locations, and annotated logical assertions over these, as expressed in Viper IL code.
The rest of this section documents and explains the most interesting of these encoding decisions, those concerning our encoding of Pancake’s machine word type, and our encoding of (shared and unshared) memory in Viper. Our objective an encoding that maximises the performance of the resulting queries to Viper’s backend, while still being sound—that is, it should never produce a Viper query that is verifiable as true, when the statement about Pancake code that it is supposed to encode is false.
We aim for most produced queries to take seconds or minutes to verify, to make verification usable as an active part of a driver developer’s workflow. Our overriding concern is to prevent the transpiler from producing queries that cause the Viper SMT-based back-end either to diverge or to take a prohibitively long time to be used for continuous integration testing. We leave proving the soundness of the encoding to future transpiler verification work (see Section 6).
3.1 Machine-Word Encoding
For the best query performance, we encode Pancake’s machine-size word variables as integers in Viper.
As a language whose only primitive type is machine-size words, Pancake’s word variables have bitvector semantics and overflows are well-defined behaviour, i.e., the variable is wrapped modulo the word size.
To preserve this semantics when encoding these as integers in Viper, which are signed and unbounded, one option is to treat arithmetic operations as modulo the word size. However, this approach significantly slows verification due to the mixing of arithmetic that is linear (such as addition, subtraction etc.) versus non-linear (such as modulo and bitvector operations) with respect to Viper’s unbounded signed integer space. Poor query performance due to mixing of different theories that require handling by different solver strategies is a well-known hazard for SMT solvers [Jovanovic and Barrett, 2013].
Instead, since overflows are rarely intended behaviour, we adopt Prusti’s approach, treating overflows as verification failures. We do this by checking the bounds of every variable after having unrolled all arithmetic operations into three-address code. We make this decision to ensure soundness at the mild expense of disallowing intended machine word overflows. These were used in the original C implementation of the driver but could be avoided with a one line change in the Pancake version.
Another concern is that bitvector operations are also non-linear with respect to Viper integers, causing similar performance concerns. Observing that almost all bitvector operations in our driver occur as part of bit masking, shifting etc., for accesses to device memory as part of a well defined device interface, we instead abstract this device interface, which we cover in Section 3.3. Additionally we precompute constant expressions and apply heuristics to rewrite common bitvector operations, e.g. x&255 is rewritten as x%256. Although the rewritten operations still use non-linear arithmetic, performance is improved compared to bitvector operations. This eliminates the vast majority of bitvector operations, with the remaining few in our Ethernet driver turning out to be reasonably performant as Viper queries.
3.2 Local Memory Encoding
As mentioned in Section 2, Pancake disallows pointers to stack variables, which simplifies modelling them in Viper: we do not have have to manage any kind of access permission or to check for references to invalid memory.
For all other memory, the memory model of Pancake poses unique challenges due to its assembly-like nature, such as the lack of first-order support for arrays and reliance on pointer arithmetic for memory operations. Due to the lack of information about what different memory regions represent—be they arrays, structs or other data structures—we must adopt the most naive modelling approach: with the exception of shared and device memory (described in the next section), we encode memory as an array of words. This approach results in non-idiomatic Viper code but captures Pancake’s native word-size treatment of memory operations accurately.
The various memory regions a function needs access to, and with which permission, are added via annotations encoded as an iterated separating conjunction [Müller et al., 2016a], which can be verified efficiently in Viper.
3.3 Shared and Device Memory Encoding
As mentioned in Section 2.4, Pancake includes dedicated load/store operations for shared memory regions, whose behaviour in a sense resembles volatile variables in C, whose accesses cannot be reordered or optimized away by the compiler. Thus, driver code uses these primitives for accesses to device memory, as well as accesses to memory shared between it and other OS components. Given this sharing, we cannot encode shared memory in Viper simply as an array of words, because we cannot rely on its contents not to change between accesses.
Device register accesses, needed for implementing drivers, make extensive use of bitwise operations to access the correct bits. As stated before, these operations result in bad performance so we seek to avoid them where possible.
2 method store_rx_free(heap: IArray,
3 device: Ref, addr: Int, value: Int)
4
5/@ shared rw u64 rx_free[lower..upper] @/
Listing 2: Method signature of shared memory store in device model and the corresponding top-level annotation in Pancake, specifying that shared memory accesses to the address range lower..upper should transpile to a Viper invocation of store_rx_free or load_rx_free.
To limit these, we model accesses to shared memory as separate Viper method calls, which the driver developer should specify in an external Viper file representing the device model for the driver’s target device. These methods define valid operations for specific address ranges corresponding to particular device registers and memory regions. They also specify requires and ensures clauses for the hardware interfaces, as assertions in terms of a global device state and the non-device memory. This is a good fit with how the Pancake semantics models shared memory operations (Section 2) as observable events, whose interpretation is parameterised on a model of the environment—a semantics which is also preserved by the Pancake compiler.
For interactions via shared memory with other OS components, we make it the responsibility of the driver developer to specify a neighbouring component model similarly in a separate Viper file. The developer should use this to capture the guarantees the driver should meet for the Viper verification to enforce, as well as any assumptions about the behaviour of those neighbouring components with respect to the shared memory.
We then provide a syntax for top-level Pancake annotations that allow the driver developer to specify the correct method(s) to use for a shared memory operation, according to the address range the driver interacts with, allowing the transpiler to infer automatically which Viper method the produced Viper model should invoke in place of the shared memory interaction (see 2).
This way, the driver developer is not required to specify the model of the device or neighbouring component before they implement their driver in Pancake. Furthermore, they do not have to modify their Pancake driver’s shared memory accesses after specifying these models—they instead add extra annotations to specify which Viper methods the transpiler should produce in place of loads/stores to given shared memory addresses. This approach ensures the separation of driver implementation and device specification, whilst also improving verification speeds.
4 Verified Ethernet Driver
4.1 Overview
We formally verify a single-core i.MX Ethernet driver for LionsOS [Heiser et al., 2025]. LionsOS uses a simple OS-side interface for device drivers with zero-copy shared-memory communication for network data. The interface employs lock-free, bounded, single-producer, single-consumer (SPSC) queues that contain meta data buffers for data addresses and lengths, as well as control information for signaling requests. The driver synchronises with the rest of the OS via semaphores (implemented as seL4 Notifications).
The target driver, implemented in Pancake, controls the MAC-NET 1 Gb/s Ethernet core common to NXP i.MX 8M Mini, Dual, QuadLite and Quad Applications Processors present in various Arm-based NXP system-on-chips (SoC), operating as a network interface card (NIC). The NIC uses DMA descriptor rings for passing the addresses of data packets. This driver will be made available for LionsOS as a choice of network drivers alongside the C version. The driver incorporates formal specifications through annotations using Hoare logic and the Viper verification framework.
Figure 3: Driver verification workflow.
Component Line Count
C Driver 351
Pancake Driver 411
Pancake annotations 558
Viper Device model 391
Viper OS interfacing model 298
Table 1: Lines of code comparisons between the Ethernet driver in Pancake implementation and C implementation, and the lines of code of our verification artifacts.
Verification workflow
Figure 3 shows the resulting verification workflow. The programmer annotates the Pancake source of the driver and processes the resulting source with the transpiler, which produces the Viper input. The programmer also supplies Viper specifications of the device interface as well as the interface between the driver and the rest of LionsOS. These are then processed by Viper, which returns a result (proved or falsified) or times out. Table 1 shows the total size of our verification artifacts.
Device model
We model the hardware NIC device and its non-deterministic state in Viper specifications. The model is then connected to Pancake driver source code after transpilation, using the top-level Pancake annotations as described in Section 3.3.
The device state consists of RX (receive) and TX (transmit) hardware descriptor rings, with each ring decomposed into three integer sequences representing data addresses, lengths, and bitfields. These integers represent machine words and are converted to bitvectors when bit-level operations are required. As discussed above, by isolating the descriptor components, such a decomposition simplifies verification and reduces SMT solver complexity, particularly when using non-linear arithmetic and bitvector operations. Each shared memory operation corresponds to a Viper method in the device model, with memory loads returning non-deterministic values within valid ranges, and memory stores satisfying both address and value requirements.
We model the non-determinism of the values the driver could possibly obtain from the device by underspecifying its interface methods, forcing the verification of the calling context of the method (i.e. the Pancake code that invoked the shared memory operation) to account for a wide range of possible values constrained only by their ensures clauses. For example, load_EIR in 3 ensures that the return value is a valid unsigned 32-bit word value, as captured by bounded32(retval). Verification must subsequently succeed for any value returned by that interface that satisfies the ensures, i.e., any unsigned 32-bit word value.
What is verified
This verification work establishes the following four classes of critical guarantees:
  1. 1.
    Compliance with NIC device interfaces (Section 4.2): we verify that:
    • device memory accesses are within the right address ranges; and
    • values written to the device obey the device’s requirements.
  2. 2.
    Compliance with neighbouring OS component interfaces (Section 4.3): we verify that the driver adheres to the network queue signalling protocols, specifically:
    • for consuming meta data buffers from network queues, the driver requests the OS wake-up signals OS if hardware rings have vacancy,
    • for providing meta data buffers to the OS, the driver notifies the OS’s semaphores if and only if (i) a signal was explicitly requested by the OS and (ii) the queue state has been changed by the driver; and
    • the driver clears signal requests after notifying the OS (to prevent double signalling)
  3. 3.
    Guaranteed data integrity across transfers (Section 4.4): we verify that there is no data loss in the driver by enforcing integrity check on packet meta data (addresses and lengths) on all pathways of meta data transfers in the driver (note that, with LionsOS [Heiser et al., 2025], the driver never needs to access or map the packet data itself, and hence never violates the integrity or confidentiality of the packet data); and
  4. 4.
    Memory safety through region isolation and restricted access controls (Section 4.5): we verify that access to TX-related states and RX-related states are correctly restricted to the parts of the driver responsible for transmitting and receiving, respectively.
4.2 Device Interface Compliance
First, we enforce that the driver only accesses the parts of device memory that comprise the NIC’s hardware interfaces for packet receipt and transmission— namely, the device’s RX (receive) and TX (transmit) hardware descriptor ring regions and other essential registers. To enforce this, recall from Section 3.3 that our transpiler supports and looks for top-level Pancake annotations that specify which addresses correspond to valid device interfaces, as illustrated in 2. Our verifier will reject any Pancake drivers that attempt to invoke a shared memory load or store operation on an address that is not covered by any such annotations.
2 method store_EIR(device: Ref, addr: Int,
3 value: Int)
4 requires addr == (REG_BASE + EIR_OFFSET)
5 requires value == IRQ_MASK
6 requires valid_device(device)
7 ensures valid_device(device)
8
9method load_EIR(device: Ref, addr: Int)
10 returns (retval: Int)
11 requires addr == (REG_BASE + EIR_OFFSET)
12 requires valid_device(device)
13 ensures bounded32(retval)
14 ensures valid_device(device)
15
16/@ shared rw u32 EIR[REG_BASE + EIR_OFFSET] @/
Listing 3: Examples of device register store/load interfaces in the device model as specified by Viper methods using requires and ensures, and corresponding top-level Pancake annotation.
Second, we verify that, whenever the driver interacts with these device interfaces, it does so in the required way not to put the device into a bad state, as specified by its documentation and captured by our device model. To enforce this, we specify two kinds of requires clauses for device interface methods, as illustrated in 3 for a representative pair of examples, the store and load methods for EIR, a particular device register:
  1. 1.
    Method-specific requirements, such as the store_EIR method’s requires of both an address and a value requirement—namely, that the address is the EIR’s and that the driver only ever writes a particular IRQ_MASK constant to it;
  2. 2.
    Device-wide invariants—in this example, captured by valid_device in store_EIR’s third requires clause. For our NIC, valid_device asserts that the state of the hardware descriptors remains valid: the bitfields are cleared and set properly according to the device’s documented specifications, data pointers are 32-bit width and byte-aligned, and data lengths are within 16-bit bounds.
The device state of our NIC device, as modelled by device and asserted valid by valid_device(device), comprises the hardware descriptor rings, with each ring decomposed into three integer sequences representing data addresses, lengths, and bitfields. These integers represent machine words and are converted by Viper to bitvectors when bit-level operations are required. Our decomposition of these descriptor components into native Viper integers thus simplifies verification and reduces SMT solver complexity, particularly when using non-linear arithmetic and bitvector operations.
Note that, while the presence of valid_device in the requires clauses of device interface methods requires the driver not to violate the validity of the device state, its presence also in the ensures clauses of all device methods specifies that we can assume the device itself will maintain that same validity invariant through all driver-device interactions.
4.3 OS Communication Protocols
Using much the same approach just described in the previous section for specifying valid device interactions, we constrain the driver’s access to shared memory regions for network SPSC queues through annotations when interfacing with LionsOS components. We also model the shared SPSC queues non-deterministically to verify the driver maintains protocol compliance—assuming the neighbouring OS component maintains it too—without assuming specific state values.
4.4 Data Integrity
To ensure reliable data transfer between the OS and the device during the translation between hardware descriptor and SPSC formats, we verify that the driver maintains data integrity by tracking packet addresses and lengths.
For example, we check that the given data address and data length are stored properly after updating the TX hardware descriptor ring, as shown in 4.
2 buffer = net_dequeue(os_tx_avail);
3update_tx_hw_ring(hw_tail, buffer);
4/@ assert(device.hw_ring_tx[hw_tail].data_addr
5 == buffer.data_addr) @/
6/@ assert(device.hw_ring_tx[hw_tail].data_len
7 == buffer.data_len) @/
Listing 4: Data integrity verification example in annotated Pancake. For brevity, unwrapping of predicates referred to by the assertions is omitted.
We establish this integrity check on all pathways of data transfers in the driver. We also verify data transfer completeness by ensuring that within the driver, the number of SPSC queue operations align with the number of hardware descriptor ring state changes, so there is no data loss in the driver.
4.5 Memory Access Control
In addition to the memory access constraints described in Sections 4.2 and 4.3, we also verify that only the parts of the driver responsible for packet transmission access any TX-related descriptor rings and SPSC queue state, and likewise that only its packet receipt paths access RX-related state.
We enforce this using Viper’s native permissions to specify access controls, in this way providing formal verification of memory safety and region isolation. In effect, our driver verification applies a separation logic-like principle by partitioning the driver’s global memory into RX and TX regions (which reflects their use in LionsOS [Heiser et al., 2025]).
5 Performance Evaluation
We now examine how our verified Pancake driver compares to the original C implementation.
Pancake driver compilation time is a matter of seconds, and the verification of the driver in full takes around 20 minutes on a typical laptop. When verified separately, the device model takes 10 minutes and the driver’s functions each take around 1 minute.
Our evaluation platform is an AVnet MaaXBoard with an NXP i.MX8MQ SoC, having four Arm Cortex A53 cores capable of a maximum of 1.5 GHz. We run our measurements at a fixed clock rate of 1 GHz to avoid the need for thermal management. [^2] Note that running at a low clock rate will magnify software overheads, and as such presents a pessimistic performance. The board has 2 GiB of RAM and the on-chip 1 Gb/s NIC specified earlier (in Section 4).
The evaluation system runs a networking client on LionsOS. The client simply receives data packets from the NIC and echoes them back. We use an external load generator that sends an adjustable load (requested throughput) to the target system, and measures the amount of data received back (received throughput) as well as the latency. On the evaluation system we also measure CPU load.
Figure 4: Performance of Ethernet Driver written in Pancake vs C, in terms of achieved throughput (Xput) and Driver CPU utilisation.
Figure 5: Performance of Ethernet Driver written in Pancake vs C, in terms of average round trip time (RTT) with standard deviation and total CPU utilisation.
Figure 4 shows the result. Like the C driver, the Pancake driver has no problems handling the requested load: the received throughput equals the requested. The Pancake version of the driver uses slightly more CPU time to handle the load than the C version, the difference is about 10%. Viewed in the context of LionsOS using less than half the CPU of Linux to handle a particular network load, the slightly increased overhead of the Pancake driver is in the noise.
Figure 5 puts this slight increase into context, by looking at latency (RTT) and the overall CPU use of the whole system. The difference in CPU usage of the drivers becomes unnoticeable. Similarly, the differences in latency is well within the variance of the results. This is in the context of LionsOS using less than half the CPU of Linux to handle a particular network load [Heiser et al., 2025]. We can summarise that the performance cost of the verified driver is in the noise.
6 Discussion
While our main results are formally verified, no proof about a real-world artifact is ever fully complete and self-contained: there will always be a trusted computing base (TCB). The TCB, in brief, is everything that we rely on for the correct operation of our system, but which is currently outside the scope of the formal verification effort. For a formal verification result, the TCB is the main threat to validity.
In Section 6.1, we discuss our TCB. Section 6.2 discusses alternative verification approaches we are pursuing that would reduce the TCB further. Finally, Section 6.3 discusses potential usability improvements to the verification front-end.
6.1 Threats to Validity
For the Pancake compiler correctness proofs, we trust: that the HOL4 theorem prover is a sound implementation of higher order logic; that the official specification of the Arm ISA is correctly implemented by the CPU [Kanabar et al., 2022]; and that an unverified linker for connecting the Pancake binary to external code such as LionsOS library routines is correct. With these as the TCB, the Pancake compiler is verified, just like the CakeML compilation passes that it leverages, to preserve the semantics of the main function of the source program.
Our verification results presented here establish certain observable properties of the driver including its behaviour after reentry. Putting these together, the verified compiler guarantees that the binary obtained by compiling this driver should preserve the verified observable properties, although we are not (yet) providing links between the two results formally, in the sense that there is no theorem yet to state this on a single, unified formalism. Future work will strengthen this by verifying the transpiler as well as extending the compiler proofs to account for reentry points.
Moreover, while we plan to verify our transpiler from annotated Pancake to Viper (Section 6.2), this is not done yet. Until then, a mistake in the transpiler could produce a Viper query that the underlying SMT-based backend can prove true, even if the property as specified at the Pancake program point via Viper annotations is false. Beyond this, we also trust the Viper verification infrastructure and SMT solvers called by its backends to discharge the Hoare logic queries specified using its input language soundly, i.e., only when true.
Like any device driver, we trust the device not to malfunction—that is, to guarantee that it meets the ensures specifications we impose on return from device interfaces in our device model. This includes trusting that the device’s initialisation process establishes a valid initial state. Verification of the device hardware and initialisation process would be needed to gain further assurance of this.
Meanwhile, we assume that the neighbouring OS components comply with the SPSC queue protocols, as mentioned in Section 4.3, and that the driver acts as the sole consumer or producer per queue. We will need to leverage Viper’s access permission system to establish more sophisticated guarantees about concurrent accesses and thread safety properties.
Finally, we trust the operating system kernel and its user-level support libraries not to crash or malfunction. LionsOS runs on the seL4 OS microkernel, which is verified not to crash [Klein et al., 2014], and relies on the seL4 Microkit support library, whose main server loop has been verified not to exhibit undefined behaviour [Paturel et al., 2023]. However, functional properties of individual seL4 system calls relied on by Microkit have not yet been verified.
6.2 Future Verification Efforts
We are continuing to implement more drivers for LionsOS in Pancake and to apply the verification method presented in this paper to those drivers firstly to produce more verified drivers but also to assess and improve the verification process. We also have various verification related future projects based on Pancake, including those we discuss below.
Transpiler verification
Proving the formal correctness of the transpiler implementation, as mentioned in Section 3, is in progress. This means that the transpiler is currently included as part of the TCB, as discussed in Section 6.1. What we aim to prove here is to establish the validity at the level of the Pancake input of the statements proved on at the level of the transpiled Viper IL code. Such a proof would complete our end-to-end verification story, by allowing us to soundly infer the correctness of our initial Pancake code from a successful verification run of the corresponding Viper IL code. This still assumes the correct operation of the Viper toolchain which our verification step relies on. Efforts exist to validate parts of this toolchain [Parthasarathy et al., 2024; Gössi, 2016].
We will do this proof using the theorem prover HOL4, which Pancake compiler is implemented. Though the current Rust implementation of the transpiler contains 7 kLOC, a significant portion of this code consists of workarounds for extracting the underlying abstract syntax from an unparsed Pancake program without interfacing directly with the existing Pancake parser. We can remove these workarounds entirely by connecting the transpilation to the Pancake parser and handling its formal syntax and the semantics directly, which will simplify the implementation considerably, making the verification easier.
Interactive theorem proving for Pancake programs
The formal semantics for Pancake and the compiler for it are developed in the HOL4 interactive theorem prover. It would be natural to verify a Pancake program entirely in HOL4. We have performed a number of small experiments in this style, verifying simple example programs such as a loop that sums the elements of an array. These experiments are an important validation of the Pancake formal semantics. The compiler correctness proof establishes that the formal semantics captures the behaviour of the binary, but that is only useful if it is possible to reason at the level of the semantics.
We have done some of these proofs by directly manipulating the Pancake semantics in HOL4. We have also constructed a Hoare-style precondition-postcondition logic for Pancake programs, which is a standard approach to verification of programs in an imperative language such as Pancake. This gives us a comparison point to our Viper-based approach. Our Hoare logic is far more user-friendly than directly working with the semantics, but far less user-friendly than using the transpiler. One advantage of the Hoare logic is that its correctness results compose directly with the compiler correctness theorem in HOL4, producing a binary correctness theorem with a minimal TCB, useful for small safety-critical programs whose correctness must be certified.
Another strength of HOL4 is its general logic, which allows us to phrase and prove properties that we could not easily express in Viper. We are interested in reasoning about concurrent scenarios. LionsOS encourages systems to be composed from smaller components (such as our Ethernet driver), most of which are internally single-threaded but execute in parallel. We think that studying a parallel protocol in HOL4 and proving correctness of each Pancake component against that protocol using Viper is an interesting direction for future work.
This composition of components will be correct if each component performs the correct I/O actions. The I/O behaviour of a program can be captured in a single value using interaction trees [Xia et al., 2020]. These are a new approach to program semantics that use an infinite coinductive tree of input and output event nodes to model a program. We have developed an alternative semantics for Pancake using interaction trees, and have experimented with proofs that programs obey an I/O protocol using this semantics. This is a promising approach to verifying, for instance, a whole network sub-system.
6.3 Transpiler usability improvements
The current workflow involves transpiling Pancake into Viper for verification. Whilst the transpiler allows for direct verification, error messages are tied to the generated Viper code and are not reflected back to the Pancake code, complicating debugging. Additionally, while many annotations are automatically inferred, some require manual specification, increasing the effort required from the programmer.
In addition to the command line tool we have built an initial development environment for verified Pancake based on Microsoft’s VS Code framework. As of yet this lacks some features like support for the separate Viper device model. The transpiler supports verifying functions individually. This could be integrated into the development environment to allow for efficient re-verification of only modified functions.
Despite these shortcomings, we expect that these tools can be combined and improved to form a cohesive and usable toolchain for writing and verifying device drivers.
7 Related Work
Among various verification efforts in the OS context, a number of prior works have investigated verification of device drivers at various programming language levels. Within this space, researchers have employed different verification approaches like model checking [Kim et al., 2008] and interactive theorem proving [Möre, 2021]. Among these efforts, [Penninckx et al., 2012] developed their verification using VeriFast [Jacobs et al., 2010], a deductive verification approach similar to ours. They also notably extended their analysis to include concurrency properties beyond our current scope. However, a common limitation across all these approaches was the significant gap between the analysed model and the actual executable code—a gap which we narrow by using Pancake’s verified compilation stack.
Others, which we detail below, have done better at closing such gaps (e.g. in some cases providing support for direct access by drivers to memory-mapped device ports, as we have), but failed to demonstrate scalability beyond the simplest serial drivers – among other reasons, by failing to include any performance evaluation of the drivers they verified.
The earliest driver verification effort for a non-trivial device we are aware of is that of Alkassar and Hillebrand [2008]; Alkassar [2009], who verified a (still simplified) ATAPI hard disk device driver in Isabelle/HOL interactive proof assistant [Nipkow et al., 2002] as part of the Verisoft project. Similar to our work, they verify their driver relative to a functional model of the memory-mapped device – in their case, based on a subset of the ATAPI standard. However, the type safety of the fragment of C they used for most of their OS limited their ability to model direct access to memory-mapped device ports directly from that language; consequently, they instead had to write and verify their driver in a MIPS-like assembly language.
Duan and Regehr [2010] presented a framework for verifying device drivers integrated with the L3 model of ARM machine code [Fox, 2003; Fox and Myreen, 2010] for HOL4 [Slind and Norrish, 2008], with a UART driver as the case study. Like Alkassar and Hillebrand [2008] they did not support reasoning about DMA, but Duan [2013] later added support in the form of Hoare triples for device memory access scenarios. The way we integrate shared memory access in Pancake with the specification of requires and ensures as Viper annotations is similar, and allows us to impose the requirements of our device model on our driver’s device memory access directly from the Pancake language. Schwarz and Dam [2014] further extend the L3 model to support device drivers with DMA. This goes further than the device model of our paper, as the Ethernet device we verify is documented not to interfere with the hardware ring indices, which are left under the control of the driver.
As part of the CertiKOS project, Chen et al. [2016] added support for verifying drivers and integrating them with their OS verification framework in the Rocq (formerly Coq) interactive proof assistant. Unlike the above works, whose drivers were implemented in assembly, the serial and interrupt controller device drivers Chen et al. [2016] verified are implemented in ClightX [Gu et al., 2015], an extension of the CompCert Clight language [Leroy, 2009] with extra instrumentation to support CertiKOS’s abstraction-guided approach to OS verification. Like our work, their use of a verified compiler (a modification of CompCert) to compile the driver down to binary gives some assurance that any properties proved at the driver source level are preserved down to the binary. However, driver verification in their framework requires interactive proofs in Rocq, for a C variant whose proof relies on adding abstract state elements that can influence program execution. This is more disruptive to the original code than mere annotations or typical “ghost state”, and arguably requires more formal methods experience than automated deductive verification via annotations.
Unlike our work, none of the works above presented any analysis or discussion their drivers’ performance.
The Ironclad verification [Hawblitzel et al., 2014] includes a network driver. They verify assembly code by translating it into Dafny code along with assertions. They argue that verification at the level of assembly avoids issues with the compiler; since we have a verified compiler we prefer to verify the sources that the programmer writes. Like in our work, the Ironclad authors annotate each driver function. They mostly establish memory safety and address a public/private data distinction. There does not appear to be any clear statement of the function of the driver or of functional correctness. The Ironclad paper gives some performance data, showing overheads of up to two orders of magnitude for single requests to their verified versus their unverified system, but no measurements of the driver alone or of the system under load, which would be needed to make a direct comparison to our work. They also note that performance is not their priority, whereas our goal here is to achieve both of high guarantees and performance.
Erbsen et al. [2021] offers an end-to-end story of a simple embedded system that includes an Ethernet driver. However, this system is simple enough that it does not require an operating system in the usual sense. They also report ten-fold overhead as the performance results for the whole system. Again, their work does not provide analysis on the driver by itself, making a direct performance comparison difficult.
We are also aware of some current efforts by Chen et al. [2024] to verify device drivers written in Rust [Klabnik and Nichols, 2017] using the Verus automated deductive verifier [Lattuada et al., 2023]. However, without a formal semantics let alone a verified compiler, the possibility of end-to-end assurances for Rust-based drivers still seems remote.
Finally, there has been also been work on driver synthesis by [Ryzhyk et al., 2009b, 2014] that took as input detailed specifications of interfaces for (1) the device class the driver needs to implement, the (2) device itself and (3) OS service it needs to provide to the rest of the OS, written in a custom specification language. Although, like in our work, their device interface included details such as valid registers and their sizes, it also included more detailed elements like a state transition diagram. In our work, we have a model of device state that we use for specifying and verifying the maintenance of invariants (the valid_device(device) assertion seen in 3 and explained in Section 4.2)—this could in future form the basis for more detailed, state machine-based specifications of internal device states. Note, however, that this synthesis work could not deal with DMA.
8 Conclusion
This paper presents, to our knowledge, the first formal verification of a demonstrably performant driver for a realistic, non-trivial device, the Ethernet NIC common to a number of variants of NXP i.MX 8M processors.
It also introduces the Pancake systems programming language, designed especially for systems-level code to be amenable to formal verification. With Pancake, it makes two enabling contributions: (1) a verified compiler that carries the semantics of Pancake down to binary, leveraging CakeML’s verified compiler backend; and (2) an automated deductive verification front-end that takes Pancake with Viper annotations, leveraging the Viper SMT-based verification framework.
This work shows that Pancake is usable for developing verified, performant drivers. A PhD student with a systems background and not much formal methods experience was able to write and verify the aforementioned Ethernet driver in a few person-months. The Pancake driver shows performance very close to C.
This work paves the way for verified development of performant device drivers—a leading source of OS vulnerabilities—as common-place infrastructure.
Acknowledgements
Many thanks to Alessandro Legnani’s project co-supervisors Toby Murray, for his role in envisioning the use of Viper, and Peter Müller for his feedback on this. We thank also Krishnan Winter, Benjamin Nott, Craig McLaughlin and Remy Sasseau for their past and ongoing contributions to other aspects of Pancake. Finally, we thank Isitha Subasinghe and Zoltan Kocsis, for their help understanding SMT performance issues from theory mixing.
The development of Pancake was made possible through the generous support of multiple organisations: the UAE Technology Innovation Institute (TII) under the Secure, High-Performance Device Virtualisation for seL4 project, DARPA under prime contract FA 8750-24-9-1000, UK’s National Cyber Security Centre (NCSC) under project NSC-1686, and the German Agentur für Innovation in der Cybersicherheit GmbH (Cyberagentur). Finally, Junming Zhao gratefully acknowledges the topup scholarship donated by the winners of the 2023 ACM Software System Award.